OpenStudy (anonymous):

When 17 is divided by k, where k is a positive integer less than 17, the remainder is 3. What is the remainder when the sum of the possible values of k is divided by 17? (a) 2 (b) 3 (c) 4 (d) 5 (e) 6

OpenStudy (istim):

Well, what do we know?

OpenStudy (anonymous):

well k is less than 17 and is positive

OpenStudy (istim):

Wait....ignore me...I'm...doing other stuff right now, sorry.

OpenStudy (istim):

Someone else, please help!

OpenStudy (zzr0ck3r):

the numbers are 7 and 14, 17/14 remainder of 3 17/(2*7) same thing 14+7 = 21 so 21/17 = 4

OpenStudy (zzr0ck3r):

remainder of 4*

OpenStudy (alexwee123):

lol this really sounds like a SAT 1 questions :p

OpenStudy (alexwee123):

question*

OpenStudy (anonymous):

it is :p

OpenStudy (zzr0ck3r):

so the answer is 4

OpenStudy (helder_edwin):

u r dividing 17 by k and have remainder 3 then u have the following equation \[ \large 17=kq+3 \]

OpenStudy (helder_edwin):